From the data in the table below, estimate two regression models using a calculator or MS Excel (calculations only) to solve this. At a very minimum, please attempt one using a calculator. Show your work in a table Model 1: Weekly Disposable IncomeAge+ Model 2: Weekly Per Capita Consumption Age+ Age (years) Income (S Weekly Disposable Weekly Consumption per capita (S 30 35 40 45 50 500 550 600 500 900 1000 1000 1200 1500 50 60 60 50 70 75 80 90 60 65 70 For each model a) Find the slope, intercept and model equation. b) Interpret the estimated regression line c) Find and interpret R2 d) Find the predicted value of the dependent variable when Age is 58.
Answer #1

for weekly disposable income-

X(age) Y(weekly disposable income) (x-x̅)² (y-ȳ)² (x-x̅)(y-ȳ)
30 500 400 130401.2 7222.222
35 550 225 96790.12 4666.667
40 600 100 68179.01 2611.111
45 500 25 130401.2 1805.556
50 900 0 1512.346 0
55 1000 25 19290.12 694.4444
60 1000 100 19290.12 1388.889
65 1200 225 114845.7 5083.333
70 1500 400 408179 12777.78
ΣX ΣY Σ(x-x̅)² Σ(y-ȳ)² Σ(x-x̅)(y-ȳ)
total sum 450 7750 1500 988888.9 36250
mean 50 861.111111 SSxx SSyy SSxy

sample size ,   n =   9          
here, x̅ =   50       ȳ =   861.1111111  
                  
SSxx =    Σ(x-x̅)² =    1500          
SSxy=   Σ(x-x̅)(y-ȳ) =   36250          
                  
slope ,    ß1 = SSxy/SSxx =   24.16666667          
                  
intercept,   ß0 = y̅-ß1* x̄ =   -347.2222222          
                  
so, regression line is   Ŷ =   -347.2222   +   24.1667   *x
                  

correlation coefficient ,    r = Sxy/√(Sx.Sy) =   0.9412          
                  
R² =    (Sxy)²/(Sx.Sy) =    0.8859  

-----------------------------

a)

slope = 24.167

intercept= -347.2222

eqn is

Ŷ =   -347.2222   +   24.1667   *x


b) for every unit increase in age, weekly disposable income will get increase by 24.1667

c) R² = 0.8859

88.59% variations in observations of variable Y(weekly disposable income) is explained by variable X(age)

d)

age,X=58

Ŷ =   -347.2222   +   24.1667   *58=1054.444

-----------------------------------------------------------------------------------------------------------------------------------------------------------------

for weekly consumption per capita-

X Y (x-x̅)² (y-ȳ)² (x-x̅)(y-ȳ)
30 50 400 241.9753 311.1111
35 55 225 111.4198 158.3333
40 60 100 30.8642 55.55556
45 60 25 30.8642 27.77778
50 50 0 241.9753 0
55 70 25 19.75309 22.22222
60 75 100 89.19753 94.44444
65 80 225 208.642 216.6667
70 90 400 597.5309 488.8889
ΣX ΣY Σ(x-x̅)² Σ(y-ȳ)² Σ(x-x̅)(y-ȳ)
total sum 450 590 1500 1572.222 1375
mean 50 65.5555556 SSxx SSyy SSxy

sample size ,   n =   9          
here, x̅ =   50       ȳ =   65.55555556  
                  
SSxx =    Σ(x-x̅)² =    1500          
SSxy=   Σ(x-x̅)(y-ȳ) =   1375          
                  
slope ,    ß1 = SSxy/SSxx =   0.916666667          
                  
intercept,   ß0 = y̅-ß1* x̄ =   19.72222222          
                  
so, regression line is   Ŷ =   19.7222   +   0.9167   *x
                     
                  
correlation coefficient ,    r = Sxy/√(Sx.Sy) =   0.8954          
                  
R² =    (Sxy)²/(Sx.Sy) =    0.8017  

------------------------------

a)

slope = 0.9167

intercept= 19.7222

eqn is

Ŷ =   19.7222   +   0.9167   *x

b) for every unit increase in age, weekly per capita consumption will get increase by 0.9167

c) R² = 0.8017

80.17% variations in observations of variable Y is explained by variable X(age)

d)

age,X=58

Ŷ =   19.7222   +   0.9167   *58 = 72.889